The addition of dimethylglioxime, H2C & H & O2N2, to a solution containing the nickel ion (I) results in a precipitate: 2H Nickel dimethylglyximex is a bulky precipitate that is inconvenient to handle 2H2C4H & O2N2 Ni (HC4H & O2N2) z + in quantities greater than 175 mg The amount of nickel in an alloy type of a magnet is 24%. Calculate the sample size that should not be exceeded when these alloys are analyzed to determine nickel
Answer #1

The complex of dimethylglyoxime (DMG) with Ni ion has the formula Ni(DMG)2. So there exists 1 molecule of Ni per molecule of the complex. The molar mass of the complex is 288.917g/mol.

Given that 175 mg is the maximum convenient quantity that can be handled, the moles of the complex in it is 175/288.917 = 0.6057 mmol (1 mmol = 0.001 mol)

So, the moles of Ni in that will also be 0.6057 mmol. This amounts to (0.6057 mmol)*(58.6934 g/mol) = 35.5505 mg.

Given that 24% of the alloy is Ni, the sample taken should be such that 24% of it is 35.5505 mg, since alloys are homogenous solid solutions and the concentration of Ni will be equal throughout the sample. This amounts to 35.5505/0.24 = 148.1271 mg of the sample.

So, a sample size of 148.1271 mg of alloy should not be exceeded to ensure that the resultant DMG complex does not exceed the limit.